# Synthetic gain for electron-beam spectroscopy

### v1

@@ Abstract
Electron-beam microscopy and spectroscopy featuring atomic-scale spatial resolution have become essential tools used daily in almost all branches of nanoscale science and technology. As a natural supercontinuum source of light, free electrons couple with phonons, plasmons, electron-hole pairs, inter- and intra-band transitions, and inner-shell ionization. The multiple excitations, intertwined with the intricate nature of nanostructured samples, present significant challenges in isolating specific spectral characteristics amidst complex experimental backgrounds. Here we introduce the approach of synthetic complex frequency waves to mitigate these challenges in free-electron–light interaction. The complex frequency waves, created through causality-informed coherent superposition of real-frequency waves induced by free electrons, offer virtual gain to offset material losses. This amplifies and enhances spectral features, as confirmed by our electron energy loss and cathodoluminescence measurements on multi-layer membranes, suspended nanoparticles, and film-coupled nanostructures. Strikingly, we reveal that our approach can retrieve resonance excitation completed buried underneath the zero-loss peak, substantially enhance the quality of hyperspectral imaging, and resolve entangled multiple-photon–electron events in their quantum interaction. Our findings indicate the versatile utility of complex frequency waves in various electron-beam spectroscopy and their promising diagnostic capabilities in free-electron quantum optics.
@@
Next, we move to cathodoluminescence, another important type of spectroscopy that combines high spatial and spectral resolution. Fig. 4a illustrates the experimental structure—a film-coupled nanoantenna geometry consisting of single silver nanoparticles of diameter on a thick gold film, interacting with a electron beam. Again, we perform a hyperspectral CL imaging measurement along a scanning line across the nanoparticle (see red line in Fig. 4b). In the raw hyperspectral image as measured (Fig. 4c), the trajectory of only a single mode can be well recognized at , together with faint features of another mode around . Nevertheless, after CFW processing, four modes appear in total, and they are labeled as M1 to M4 with ascending frequency as shown in Fig. 4d. The spectra at representative positions A, B, and C before and after CFW processing are shown in Fig. 4e and f, respectively. To elucidate the origin of each mode, we perform numerical simulations, and the resulting spectra before and after CFW processing are shown in Fig. 4g, which corresponds well with the experimental data in Fig. 4e and f. We further elucidate the origin of the peaks using eigenmode calculations in Fig. 4h. Because of the sphere-film–coupled geometry, the gap plasmon modes and the Mie modes get hybridized. In the low-frequency region, M1 and M2 predominately display gap-plasmon features and are labeled by their quantized in-plane wavevectors and along the radial and azimuthal direction, respectively Yang et al. (2017). As the frequency increases, M3 and M4 primarily exhibit Mie characteristics and can be identified by the integer labels of spherical harmonics. Nevertheless, CL in this region becomes highly multi-channel in . Therefore, in Fig. 4, we display the eigen-fields of and for M3 and M4, respectively. These modes are the primary contributors to these two peaks according to the Mie CL calculation De Abajo (2010); De Abajo and Howie (1998).
@@
Acknowledgments. We thank Germaine Arend, Frankie Y. F. Chan, Mathieu Kociak, Zemeng Lin, and Bengy T. T. Wong for experimental help and stimulating discussions.

### v2

@@ Abstract
Electron-beam microscopy and spectroscopy featuring atomic-scale spatial resolution have become essential tools used daily in almost all branches of nanoscale science and technology. As a natural supercontinuum source of light, free electrons couple with phonons, plasmons, electron-hole pairs, inter- and intra-band transitions, and inner-shell ionization. The multiple excitations, intertwined with the intricate nature of nanostructured samples, present significant challenges in isolating specific spectral characteristics amidst complex experimental backgrounds. Here we introduce the approach of synthetic complex frequency waves to mitigate these challenges in free-electron–light interaction. The complex frequency waves, created through causality-informed coherent superposition of real-frequency waves induced by free electrons, offer virtual gain to offset material losses. This amplifies and enhances spectral features, as confirmed by our electron energy loss and cathodoluminescence measurements on multi-layer membranes, suspended nanoparticles, and film-coupled nanostructures. Strikingly, we reveal that our approach can retrieve resonance excitation completely buried underneath the zero-loss peak, substantially enhance the quality of hyperspectral imaging, and resolve entangled multiple-photon–electron events in their quantum interaction. Our findings indicate the versatile utility of complex frequency waves in various electron-beam spectroscopy and their promising diagnostic capabilities in free-electron quantum optics.
@@
Next, we move to cathodoluminescence, another important type of spectroscopy that combines high spatial and spectral resolution. Fig. 4a illustrates the experimental structure—a film-coupled nanoantenna geometry consisting of single silver nanoparticles of diameter on a thick gold film, interacting with a electron beam. Again, we perform a hyperspectral CL imaging measurement along a scanning line across the nanoparticle (see black line in Fig. 4b). In the raw hyperspectral image as measured (Fig. 4c), the trajectory of only a single mode can be well recognized at , together with faint features of another mode around . Nevertheless, after CFW processing, four modes appear in total, and they are labeled as M1 to M4 with ascending frequency as shown in Fig. 4d. The spectra at representative positions A, B, and C before and after CFW processing are shown in Fig. 4e and f, respectively. To elucidate the origin of each mode, we perform numerical simulations, and the resulting spectrum is shown in Fig. 4g, which corresponds well with the experimental data in Fig. 4f. We further elucidate the origin of the peaks using eigenmode calculations in Fig. 4h. Because of the sphere-film–coupled geometry, the gap plasmon modes and the Mie modes get hybridized. In the low-frequency region, M1 and M2 predominately display gap-plasmon features and are labeled by their quantized in-plane wavevectors and along the radial and azimuthal direction, respectively Yang et al. (2017). As the frequency increases, M3 and M4 primarily exhibit Mie characteristics and can be identified by the integer labels of spherical harmonics. Nevertheless, CL in this region becomes highly multi-channel in . Therefore, in Fig. 4, we display the eigen-fields of and for M3 and M4, respectively. These modes are the primary contributors to these two peaks according to the Mie CL calculation De Abajo (2010); De Abajo and Howie (1998).
@@
Acknowledgments. We thank Germaine Arend, Frankie Y. F. Chan, Anthony H. W, Choi, Mathieu Kociak, Zemeng Lin, and Bengy T. T. Wong for experimental help and stimulating discussions.